Opened 9 years ago

Closed 9 years ago

Last modified 3 years ago

#345 closed flight processing (fixed)

EU10/01, 178/2010, Tharandt

Reported by: jb362 Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data Location: /users/rsg/airborne/workspace/EU10_01-2010_178_Tharandt

This project has not been fastcopied into the repository.

Data arrived from ARSF via network transfer on 05/07/2010.

Scientific objective: 1.Key impact factors on CO2 and H2O fluxes
1.1 examine the influence of within-site elevation changes using flux tower observations, airborne LiDAR data, and ecosystem models data (Loobos)
1.2 develop an approach to allocate CO2 and H2O fluxes to specific tree species of a mixed forest (Norunda)
1.3 assess the influence of forest thinning on CO2 uptake (Norunda)
1.4 evaluate above findings and test transferability of findings (Tharandt)
1.5 investigate the impact of storm damage (Tharandt)

  1. Reducing uncertainty in the upscaling of greenhouse gases

2.1. develop an approach to evaluate airborne flux measurements using flux tower data, footprint model and airborne LiDAR data (Loobos)
2.2 use airborne LiDAR data to validate GLAS/ICESat for different forest biomes, where accurate tree height and vegetation cover density will be co-located with GLAS footprints
2.3 evaluate the ability of GLAS spatial sampling density to represent the wider field site
2.4 develop an approach to model CO2 and H2O fluxes on a regional scale using GLAS/ICESat data and generate key driving parameters for land-surface models
2.5 evaluate whether regional CO2/H2O flux models driven by satellite optical data can be improved by the addition of tree height information from satellite LiDAR

Priority: a5

PI: N. Klujn
Sensors: ΒΆ

  • Eagle (26/08/2010)
  • Hawk (21/12/2010)
  • Leica LIDAR (26/08/2010)
  • Photography (27/07/2010)

Change History (35)

comment:1 Changed 9 years ago by jb362

Base station data is now available for this project

comment:2 Changed 9 years ago by ella

This project has not yet been fast copied into the arsf repoistry.
IPAS navigation processing done.
Applanix navigation processing done.

comment:3 Changed 9 years ago by emca

  • Owner set to emca
  • Status changed from new to accepted

comment:4 Changed 9 years ago by emca

Starting LIDAR processing

comment:5 Changed 9 years ago by ella

Hyperspectral processing done and ready for checking at ~arsf/workspace/EU10_01-2010_178_Tharandt/delivery/20100726/EU10_01.

On the flight logsheet there were three lines named 'X2' but have been change to X2a X2b and X2c.

comment:6 Changed 9 years ago by emca

Photograph processing complete. Removed the following photographs from
the delivery as they were over/under exposed or not relevant to the project
001-055, 059, 096, 131, 155, 174, 191, 192, 207, 211, 222, 229, 230,
235, 246, 252, 267, 376, 380, 385, 389, 392

Ready for delivery check at
~airborne/workspace/EU10_01-2010_178_Tharandt/delivery/20100723/EU10_01

comment:7 Changed 9 years ago by ella

Starting the check for the RCD.

[Have removed 'camera' from the Delivery Checking/Processing Order wiki page for this project to avoid it being checked twice.]

comment:8 Changed 9 years ago by ella

Camera delivery check complete. All checks out apart from:

  • 5 Readmes - delete all apart form pdf
  • hidden .svn directory in doc directory - removed

comment:9 Changed 9 years ago by jb362

Starting hyperspectral delivery check

comment:10 Changed 9 years ago by emca

  • Description modified (diff)

Photography data sent to PI 29/07/2010

comment:11 Changed 9 years ago by jb362

Hyperspectral check complete. Everything fine although:

1) Pixel overflow for lev1 images in bands:

e1: 0-122
e2: 0-98
e3: 0-80
e4: 0-97
e5: 0-99
e6: 0-114
e7: No pixel overflow
e8: 0-121
e9: 0-98
e10: 0-135
e11: 0-121
e12: 0-122
e13: 0-147
e14: 0-152
e15: 0-152
e16: 0-149
e17: 0-152
e18: 0-136
e19: 0-135
e20: 0-138
e21: 0-135
e22: 0-151
e23: 0-133
e24: 0-118
e25: 0-135
e26: No pixel overflow
e27: 0-136
e28: 0-136
e29: 0-69
e30: 0-129
e31: 0-129
e32: 0-120
e33: 0-129
e34: Header file size error

h1: 81-116
h2: No pixel overflow
h3: No pixel overflow
h4: Band 99 only
h5: No pixel overflow
h6: 82-112
h7: No pixel overflow
h8: 82-110
h9: No pixel overflow
h10: 44-51, 78-112
h11: 82-113
h12: 46-47, 80-117
h13: 44-51, 78-123
h14: 38-53, 78-124
h15: 36-53, 78-124
h16: 43-52, 78-124
h17: 37-53, 78-124
h18: 45-49, 79-122
h19: 45-48, 79-121
h20: 44-51, 78-122
h21: 45-49, 79-121
h22: 80-116
h23: 80-117
h24: 82-113
h25: 46-50, 79-120
h26: No pixel overflow
e27: 44-51, 78-122
e28: 44-51, 78-122
e29: No pixel overflow
e30: 79-116
e31: 80-115
e32: 82-113
e33: 79-117
e34: No pixel overflow

2) Band 160 blank in all lev1 images

3) Capital letter needed in comments on data quality: "Some cloud patches on a number of flight lines"

4) Move Read_Me-20100726.tex file from delivery directory

comment:12 Changed 9 years ago by emca

LiDAR processing complete and ready for delivery check at:
~airborne/workspace/EU10_01-2010_178_Tharandt/delivery/20100818

comment:13 Changed 9 years ago by ella

Starting LiDAR delivery check.

comment:14 Changed 9 years ago by ella

LiDAR delivery check complete...everything looks good.

Eagle hyperspectral delivery at:
/users/rsg/arsf/workspace/EU10_01-2010_178_Tharandt/delivery/20100726/EU10_01

(Hawk level 1 files, screenshots and mosaics are moved to /users/rsg/arsf/workspace/EU10_01-2010_178_Tharandt/delivery/hawk_20100726)

comment:15 Changed 9 years ago by emca

LiDAR and Eagle data dispatched to PI 26/08/2010 along with Photography data from day 168 (ticket ​#341)

Hawk data has not yet been sent due to calibration issue

comment:16 Changed 9 years ago by emca

LAS files made available to PI 27/10/2010 via FTP

comment:17 Changed 9 years ago by emca

Hawk images reprocessed - some of the SCT values do not look correct (lines 01, 11, 15).
Not creating delivery yet- this project will need to be looked at again.

comment:18 Changed 9 years ago by emca

  • Description modified (diff)

Reprocessed the Hawk lines, and updated some of the Hawk SCT values. Hawk delivery has been created. Eagle and Hawk SCT values are as follows:

Line Eagle Hawk
1 0.01 0.12
2 0.00 0.04
3 0.01 0.01
4 0.02 0.05
5 0.00 -0.02
6 0.01 0.00
7 0.00 0.02
8 0.02 0.04
9 -0.01 -0.02
10 -0.01 0.00
11 -0.08 0.71
12 0.01 0.00
13 0.01 -0.01
14 0.02 0.09
15 0.00 0.01
16 0.01 0.02
17 0.00 0.00
18 0.00 0.03
19 0.00 -0.01
20 0.02 0.04
21 0.01 0.00
22 0.01 -0.06
23 0.00 0.04
24 0.01 -0.01
25 -0.01 0.02
26 0.02 0.03
27 0.01 -0.01
28 0.00 0.04
29 0.03 -0.01
30 0.00 0.04
31 0.04 0.00
32 0.03 0.01
33 0.02 -0.01
34 0.01 0.01

comment:19 Changed 9 years ago by iopa

Delivery check complete,

data and delivery are in order.

Perhaps the cloud cover should be mentioned in the readme
and that it led to areas of overflow on multiple flightlines.

Otherwise good to go.

comment:20 Changed 9 years ago by emca

Hawk Lev1 one files reprocessed and copied to delivery directory

comment:21 Changed 9 years ago by iopa

Hawk has been re-checked, all seems in order, ready to re-deliver.

comment:22 Changed 9 years ago by emca

  • Description modified (diff)

Hawk data dispatched 21/12/2010

comment:23 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa
  • Status changed from accepted to assigned

This project has been r-synced.

Beginning archiving.

comment:24 Changed 9 years ago by knpa

This project has some FW lidar data.

comment:25 Changed 9 years ago by knpa

This project has some Tabi data.

comment:26 Changed 9 years ago by knpa

eagle lev3's missing

comment:27 Changed 9 years ago by knpa

There is webcam data for this project.

comment:28 Changed 9 years ago by knpa

Tarball created, made available to NEODC.

comment:29 Changed 9 years ago by knpa

Archived and backed up by NEODC.
Moved to archive disk.

comment:30 Changed 9 years ago by knpa

  • Milestone set to 2010 data processing completion
  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:31 Changed 7 years ago by anhi

Reprocessed hyperspectral delivery check:

  • Logsheet has some flightlines named differently to the ReadMe
  • Mention some dropped scans

comment:32 Changed 7 years ago by adbe

Changed ReadMe names and mentioned dropped scans

comment:33 Changed 7 years ago by adbe

Reprocessed hyperspectral delivered to Natascha Kljun on 10/09/2012 on disk 130

comment:34 Changed 3 years ago by dac

Archiving

Processed data available from NEODC (​http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EU10_01/EU10_01-2010_178_Tharandt). Started uploading raw data.

comment:35 Changed 3 years ago by dac

Archiving

Raw data also available from NEODC.

Note: See TracTickets for help on using tickets.